OpenStudy (anonymous):

Bit Confused on how to work out a geometry problem. So I have a problem where I have to find the degree of angle A. I'm given that Angle C is a right angle, and the length of the hypotenuse and opposite angle. Is there any formula I could use that would help me solve this because I'm just stuck on this question. Thanks to anyone that helps.

OpenStudy (vishweshshrimali5):

Yep you can use trigonometry

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):

Can you mark the things you are given in the above figure?

OpenStudy (anonymous):

 Thats what I was given. I also think I'm supposed to divide the opposite angle by the hypotenuse to get the sine, but I'm not sure what to do after that.

OpenStudy (vishweshshrimali5):

Hmm good work so far! see: \[\sin x = \cfrac{AC}{AB}\]

OpenStudy (vishweshshrimali5):

Now, you know the value of AC and AB, so you can calculate sin x Then using a calculator you can find out x! (Hint: x = arcsin (AC/AB))

OpenStudy (anonymous):

So would that be 12/31 or .38709... I just started working on these problems today so I'm struggling a bit.

OpenStudy (vishweshshrimali5):

Good work!

OpenStudy (vishweshshrimali5):

Now, using a calculator: x = arcsin (0.38709) = 67.226 degrees

OpenStudy (vishweshshrimali5):

Any doubt ?

OpenStudy (vishweshshrimali5):

Now, use the fact that: Sum of all angles of a triangle = 180 deg You know B and C, so can you find out A ?

OpenStudy (anonymous):

A little, I'm a bit confused on what an arcsine is.

OpenStudy (vishweshshrimali5):

It basically is \(\sin^{-1}\) For example: sin (30 deg) = 1/2 = 0.5 => 30 deg = \(\sin^{-1} (0.5)\) or arcsin (0.5)

OpenStudy (anonymous):

Ok I think i'm starting to get it. Also to what you said above, angle A would be 22.774 degrees.

OpenStudy (vishweshshrimali5):

Great work!!!! You are perfectly correct!!

OpenStudy (anonymous):

Thank you! I've spent the past hour going through online resources and haven't been able to understand it.

OpenStudy (vishweshshrimali5):

your welcome :)